# Kwai-STaR: Transform LLMs into State-Transition Reasoners

### v1

@@ 1 Introduction
Kwai-STaR takes three steps to incorporate state transition with LLM reasoning: (1) State space definition: We develop an action set and constrain the LLM to solve problems as a State-Transition Reasoner (STaR): The LLM takes one action from the action set at a time to transition the current state into a new state. (2) State-transition data construction. Although advanced LLMs possess remarkable instruction-following capabilities (Table 2), we still take appropriate training to help general language models master state-by-state reasoning. Hence, we construct a small-scale state-transition dataset with meticulous generation instructions and pipeline. (3) Curricular Training Strategy. Our dataset contains two types of instances: a majority of correct cases and a minority of wrong-then-verified cases from the data generator and trained reasoner. To maximize learning efficiency, our training strategy consists of two stages: a fundamental stage and an advanced stage. The former trains the model with the majority right cases, enabling it to grasp the state-transition manner and to solve relatively simple problems. The latter leverages pairs of wrong and verified cases to further strengthen the model’s proficiency.
@@ 2.2 Kwai-STaR State Space
Model CoT (G/M) Kwai-STaR (G/M)
GPT-4o 91.20/60.30 94.31/68.80
LLaMA-3.1-Inst-70B 95.10/68.00 96.04/84.60
Table 2: Accuracy of models prompted by CoT and Kwai-STaR instruction on GSM8K(G) and MATH(M) Benchmark.
@@ 2.3 State-Transition Data Construction
We initially validate the effectiveness of our state space in enhancing LLMs’ reasoning capabilities by prompt engineering: We compare the accuracy of models prompted by CoT and Kwai-STaR instructions on GSM8K and MATH benchmark. As shown in Table 2, Kwai-STaR instruction significantly improve the accuracy of GPT-4o and LLaMA-3.1. This indicate that our state space contributes the reasoning process of LLMs.
Although the pre-experiment has shown that state-of-art LLMs can follow instructions to perform state transition, the ability of smaller and weaker models may be not guaranteed. Thus, to improve reasoning ability of various LLMs broadly, we construct state-transition reasoning data with small scale and high quality (Table 3) to teach LLMs how to perform state transition.

### v2

@@ 1 Introduction
Kwai-STaR takes three steps to incorporate state transition with LLM reasoning: (1) State space definition: We develop an action set and constrain the LLM to solve problems as a State-Transition Reasoner (STaR): The LLM takes one action from the action set at a time to transition the current state into a new state. (2) State-transition data construction. Although advanced LLMs possess remarkable instruction-following capabilities, we still take appropriate training to help general language models master state-by-state reasoning. Hence, we construct a small-scale state-transition dataset with meticulous generation instructions and pipeline. (3) Curricular Training Strategy. Our dataset contains two types of instances: a majority of correct cases and a minority of wrong-then-verified cases from the data generator and trained reasoner. To maximize learning efficiency, our training strategy consists of two stages: a fundamental stage and an advanced stage. The former trains the model with the majority right cases, enabling it to grasp the state-transition manner and to solve relatively simple problems. The latter leverages pairs of wrong and verified cases to further strengthen the model’s proficiency.
@@ 2.3 State-Transition Data Construction
Although state-of-art LLMs can follow instructions to perform state transition, the ability of smaller and weaker models may be not guaranteed. Thus, to improve reasoning ability of various LLMs broadly, we construct state-transition reasoning data with small scale and high quality (Table 2) to teach LLMs how to perform state transition.